Moskowskaja, 267, Brest 224017 Belarus Brest State Technical University Rector: Petr Poita

10 April 2006 Deputy Ambassador Mr. Takayuki KOIKE

BSTU is going to held the ICNNAI-2006 conference and asking Japanese Embassy to be one of the co-sponsors. If allowed, then we guarantee that we are following the below. If the bellow are not followed, you may cancell this co-sponsorship.

1. We are responsible for everything concerning the conference;

2. We do not make any political activity and any other events which are not the objectives of this conference.